13. Find payment information on a loan (Unit 4D exercises 15-24) SA. You will be provided with a
principal, and loan terms (APR and number of years). Determine the monthly payment, total
amount paid and the percentage paid towards interest.
You want to buy a $12,000 car with no money down at a rate of 3% APR for 5 years. What is
your monthly payment, total amount paid, and percentage of that total that goes to interest?
Answer #1

Calculating Monthly Payment,

using TVM Calculation,

PMT = [PV = 12,000, FV = 0, N = 60, I = 0.03/12]

PMT = $215.62

Monthly Payment = $215.62

Total Amount Paid = 60(215.62) = $12,937.20

Interest paid = 12,937.20 - 12,000

Interest Paid = $937.20